# Efficient Multi-agent Navigation with Lightweight DRL Policy

### v1

@@ 3.1 Notation of reinforcement learning
This work models the multi-agent navigation process as POMDP. The POMDP can be defined as $<\mathcal{I},\mathcal{S},\mathcal{O},\mathcal{A},\mathcal{R},\mathcal{T},done,\gamma>$, where =$\mathcal{I}=\{1,...,N\}$ denotes $N$ agents, $\mathcal{S}$ denotes the set of states of all agents, $O=\{\mathcal{O}_{1},...,\mathcal{O}_{N}\}$ denotes the set of observable state of agents, $\mathcal{A}$ denotes the set of actions, $\mathcal{R}$ denotes the reward function, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\mapsto\mathcal{S}$ denotes the transfer function and $done$ indicates whether an episode has terminated in time or collided with an obstacle, and $\gamma\in[0,1)$ denotes the discount factor. At each time step $t$, the agent’s observation is $o^{t}=[o_{s}^{t},o_{g}^{t},o_{v}^{t}]$, where $o_{s}^{t}\in\mathbb{R}^{5\times L_{n}}$ is formed by concatenating 5 consecutive frames of LiDAR data, with $L_{n}$ being the measurement dimension of the LiDAR data; $o_{g}^{t}\in\mathbb{R}^{2}$ consists of the distance from the agent’s current position to the goal point, and the angle between the agent’s orientation and the direction towards the goal point; $o_{v}^{t}=[v_{i}^{t},\omega_{i}^{t}]\in\mathbb{R}^{2}$ consists of the agent’s linear and angular velocities. The agent $i$ receives $o_{i}^{t}$ and executes action $a_{i}^{t}$ according to policy $\pi_{\theta}(a^{t}_{i}|o^{t}_{i})$, where $\theta$ is the policy parameter. The policy $\pi_{\theta}(a^{t}_{i}|o^{t}_{i})$ is a Gaussian distribution, with the mean calculated by a policy network. The action $a^{t}_{i}=[v^{t}_{i},\omega^{t}_{i}]$ is sampled from the policy $\pi_{\theta}(a^{t}_{i}|s^{t}_{i})$. The state $s_{t}$ takes the action set ${a_{i}^{t},...,a_{N}^{t}}$ to transfer to next state $s_{t+1}$ according to $\mathcal{T}(s_{t+1}|s_{t},a_{i}^{t},...,a_{N}^{t})$. Also, the next observable state set is updated as $\mathcal{O}=\mathcal{T}(\{o_{1},...o_{N}\}|s_{t},a_{i}^{t},...,a_{N}^{t})$, and the reward set is obtained by =$\mathcal{R}=\{r(s_{1}^{t},a_{1}^{t}),...,r(s_{N}^{t},a_{N}^{t})\}$. A trajectory can be represented as $<o^{1},a^{1},r^{1},o^{2},a^{2},r^{2},...,s^{T},a^{T},r^{T}>$, where $T$ represents the number of interaction steps. Each agent $i$ aims to maximize the discounted future reward: $G_{i}^{t}=\sum_{t=0}^{T}\gamma^{t}r_{i}^{t}$. The state-value function is defined as $V^{\pi}(s)=\mathbb{E}[\sum_{t=0}^{T}\gamma^{t}R^{t}|s^{1}=s]$, and the action-value function is given by $Q^{\pi}(s,a)=\mathbb{E}[R(s,a)+\gamma V^{\pi}(s)|s^{1}=s,a^{1}=a]$. A action is chosen as $a={\rm arg\ max}_{a}Q(s,a)$.
3.2 Reward function
@@
$\begin{array}[]{l}R=r_{g}+r_{c}+r_{other},\\ r_{g}=\left\{\begin{array}[]{ll}r_{{arrival}}\qquad\qquad\quad\ \ \text{if }\left\|p_{i}^{t}-g_{i}\right\|<0.1,\\ w_{g}\left(\left\|p_{i}^{t-1}-g_{i}\right\|-\left\|p_{i}^{t}-g_{i}\right\|\right)\quad\text{otherwise,}\end{array}\right.\\ r_{c}=\left\{\begin{array}[]{ll}r_{{collision}}\qquad\qquad\quad\text{if }d_{i}^{t}<0,\\ w_{c}(0.1-d_{i}^{t}/2)\quad\text{ if }d_{i}^{t}<0.2,\\ 0\qquad\qquad\qquad\quad\text{ otherwise,}\end{array}\right.\\ \end{array}$ (1)
where $r_{other}$ is the other reward that varies in different approaches, $r_{g}$ means the distance reward, which is computed as the difference in the distance between the robot and the goal point at the current time step and the previous time step, $r_{c}$ means the obstacle penalty, $w_{g}$ and $w_{c}$ are the weight of $r_{g}$ and $r_{c}$ respectively, $d_{i}^{t}$ means the smallest distance between agent $i$ and obstacles at time step $t$. However, this reward function setup may make agents learn conservative policy. Additionally, many works impose control constraints, such as velocity and angular velocity, on $r_{other}$ to maintain action coherence. Too many constraints may lead to instability in the training process and difficulty adjusting the weights between various terms. Moreover, in some scenarios, the reward function does not adequately indicate optimal actions.
@@
Figure 1: Cases that are difficult to handle with the classical reward function.
@@
Figure 3: The overall framework diagram of our approaches is presented. The approaches can be divided into trajectory collection and policy update. During trajectory collection, each agent collects the trajectory independently and stores it in a replay buffer. The simulation platform will return each agent’s reward and observation after an action is executed. After a specific number of iterations, the trajectories stored in the replay buffer will be used to calculate the loss, and the network will then be updated based on the loss through gradient descent. The two stages alternate until a specific number of episodes is reached or the policy converges.
As shown in Figure 1, when an agent needs to navigate through the gap between two obstacles, it may incur penalties due to $r_{c}$ when approaching them, even though traversing the gap is the optimal action. Consequently, the agent may choose to circumnavigate around the obstacles instead. Besides, when multiple agents approach each other, they may only focus on the closest agent or obstacle to itself, neglecting potentially threatening agents that are further away but pose a real danger, which leads to unsafe navigation. To address the issues above, we have introduced modifications to the $r_{c}$ component, specifically by:
$\begin{array}[]{l}r^{\prime}_{c}=\left\{\begin{array}[]{ll}r_{collision}&\text{if }d_{i}^{t}<0,\\ w^{\prime}_{c}\left(l-sl_{i}^{t}\right)+r_{c}&\text{ otherwise,}\end{array}\right.\end{array}$ (2)
where $r_{arrival}=5$, $r_{collision}=-5$, $w_{g}=5$, $w^{\prime}_{c}=-0.25$, $w_{c}=-0.03$, $l$ represents the maximum length of the LiDAR rays, $sl_{i}^{t}$ represents the LiDAR information of agent $i$ at time step $t$, $w^{\prime}_{c}$ is a weight vector that varies with angular velocity, with the weight decreasing in $\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(x-\omega\Delta t)^{2}}{2\sigma^{2}}}$ fashion from $\omega\Delta t$ towards both sides, where $\sigma=0.2$, $\Delta t$ is the time interval. The schematic diagram of $w^{\prime}_{c}$ is shown in Figure 2.
@@
Using $r^{\prime}_{c}$, the policy pays more attention to the direction of the current angular velocity while maintaining the closest distance to the surrounding obstacles. This allows the policy to better cope with the mentioned situations and allows the agent to pay more attention to the choice of angular velocity.
4 Method
In this section, we will introduce the architecture of the policy network, the simulation platform, and the training process with training scenarios. We utilize the hidden state of LSTM to compensate for the lack of information in the observations and propose an attention structure to improve the policy ability in the multi-agent scenario. Besides, we propose a multi-agent simulation platform to bridge the gap between simulation and the real world. Additionally, the multi-agent training process is based on Proximal Policy Optimization (PPO, Schulman et al. (2017)), with modifications in trajectory collection. Finally, we use two scenarios for training: single-agent and multi-agent.
4.1 Network structure
We note that the agent only requires more biased actions toward the goal point when approaching it. In other situations, the agent merely needs to move in a general direction. As such, we clip the distance between the agent and the goal point to enable the policy to better distinguish between these two situations rather than directly using the distance as an input to the network.
Algorithm 1 PPO For Multi-Agent training
1:procedure PPO(environment, hyperparameters)
2:     Initialize policy network $\pi_{\theta}$, value function network $V_{\phi}$, and set hyper-parameters
3:     for $iteration=1,2,...,$ do
4:         for $t=1,2,...,T_{max}$ do
5:              for robot $i=1,2,...$ do
6:                  Run policy network $\pi_{\theta}$, collecting $\{o_{i}^{t},a_{i}^{t},r_{i}^{t}\}$.
7:                  Run value function network, collecting $\{V_{\phi}(o_{i}^{t})\}$
8:              end for
9:         end for
10:         for robot $i=1,2,...$ do
11:              for $t=T_{max},T_{max-1},...,1$ do
12:                  Calculate TD-error: $\delta_{i}^{t}=r_{i}^{t}+\gamma V_{\phi}(o_{i}^{t+1})-V_{\phi}(o_{i}^{t})$
13:                  Estimate advantages using GAE: $A_{i}^{t}=\sum_{t}^{T_{max}}(\gamma\lambda)^{t}\delta_{i}^{t}$
14:                  Calculate $V_{i}^{t}=A_{i}^{t}+V_{\phi}(o_{i}^{t})$
15:              end for
16:         end for
17:         Merge and shuffle the collected trajectories
18:         $\theta_{old}=\theta,\ \phi_{old}=\phi$
19:         for epoch in Epoch do
20:              $L^{{PPO}}=-\mathbb{E}\left(\sum_{i=1}^{N}\sum_{t=1}^{T_{max}}\text{min}\left(\frac{\pi_{\theta}(a_{i}^{t}|o_{i}^{t})}{\pi_{\theta_{{old}}}(a_{i}^{t}|o_{i}^{t})}A_{i}^{t},\text{clip}\left(\frac{\pi_{\theta}(a_{i}^{t}|o_{i}^{t})}{\pi_{\theta_{{old}}}(a_{i}^{t}|o_{i}^{t})},1-\epsilon,1+\epsilon\right)A_{i}^{t}\right)\right)$
21:              $L^{V}=\text{max}\left(\mathbb{E}\left(\sum_{i=1}^{N}\sum_{t=1}^{T_{max}}\left((V_{\phi}(o^{t}_{i})-V_{i}^{t})^{2},\left(\text{clip}\left(V_{\phi}(o_{i}^{t}),V_{\phi_{old}}(o_{i}^{t})-\epsilon,V_{\phi_{old}}(o_{i}^{t})+\epsilon\right)-V_{i}^{t}\right)^{2}\right)\right)\right)$
22:              $L^{{Entropy}}=H(\pi_{\theta}(a^{t}|o^{t}))$
23:              $L^{{Total}}=L^{{PPO}}+\alpha L^{V}+\beta L^{{Entopy}}$
24:              Update the parameters of $\pi_{\theta}$ and $V_{\phi}$ using Adam Optimizer w.r.t $L^{Total}$
25:         end for
26:     end for
27:end procedure

### v3

@@ III-A Notations and Problem Statement
This work models the multi-agent navigation process as POMDP. The POMDP can be defined as $<\mathcal{I},\mathcal{S},\mathcal{O},\mathcal{A},\mathcal{R},\mathcal{T},done,\gamma>$, where =$\mathcal{I}=\{1,...,N\}$ denotes $N$ agents, $\mathcal{S}$ denotes the set of states of all agents, $\mathcal{O}=\{\mathcal{O}_{1},...,\mathcal{O}_{N}\}$ denotes the set of observable state of agents, $\mathcal{A}$ denotes the set of actions, $\mathcal{R}$ denotes the reward function, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\mapsto\mathcal{S}$ denotes the transfer function and $done$ indicates whether an episode has terminated in time or collided with an obstacle, and $\gamma\in[0,1)$ denotes the discount factor. At each time step $t$, the agent’s observation is $o^{t}=[o_{s}^{t},o_{g}^{t},o_{v}^{t}]$, where $o_{s}^{t}\in\mathbb{R}^{5\times L_{n}}$ is formed by concatenating 5 consecutive frames of LiDAR data, with $L_{n}$ being the measurement dimension of the LiDAR data; $o_{g}^{t}\in\mathbb{R}^{2}$ consists of the distance from the agent’s current position to the goal point, and the angle between the agent’s orientation and the direction towards the goal point; $o_{v}^{t}=[v_{i}^{t},\omega_{i}^{t}]\in\mathbb{R}^{2}$ consists of the agent’s linear and angular velocities. The agent $i$ receives $o_{i}^{t}$ and executes action $a_{i}^{t}$ according to policy $\pi_{\theta}(a^{t}_{i}|o^{t}_{i})$, where $\theta$ is the policy parameter. The policy $\pi_{\theta}(a^{t}_{i}|o^{t}_{i})$ is a Gaussian distribution, with the mean calculated by a policy network. The action $a^{t}_{i}=[v^{t}_{i},\omega^{t}_{i}]$ is sampled from the policy $\pi_{\theta}(a^{t}_{i}|s^{t}_{i})$. The state $s_{t}$ takes the action set ${a_{i}^{t},...,a_{N}^{t}}$ to transfer to next state $s_{i}^{t+1}$ according to $\mathcal{T}(\mathcal{S}^{t+1}|\mathcal{S}^{t},a_{i}^{t},...,a_{N}^{t})$. Also, the next observable state set is updated as $\mathcal{T}(O^{t+1}|\mathcal{O}^{t},a_{i}^{t},...,a_{N}^{t})$, and the reward set is obtained by =$\mathcal{R}=\{r(s_{1}^{t},a_{1}^{t}),...,r(s_{N}^{t},a_{N}^{t})\}$. A trajectory of agent $i$ can be represented as $<o_{i}^{1},a_{i}^{1},r_{i}^{1},o_{i}^{2},a_{i}^{2},r_{i}^{2},...,s_{i}^{T},a_{i}^{T},r_{i}^{T}>$, where $T$ represents the number of interaction steps. Each agent $i$ aims to maximize the discounted future reward: $G_{i}^{T}=\sum_{t=0}^{T}\gamma^{t}r(s_{i}^{t},a_{i}^{t})$. The state-value function is defined as $V^{\pi}(s_{i}^{T})=\mathbb{E}[G_{i}^{T}|s^{1}_{i}]$, and the action-value function is given by $Q^{\pi}(s_{i}^{T},a_{i}^{T})=\mathbb{E}[G_{i}^{T}+\gamma V^{\pi}(s_{i}^{T})|s^{1}_{i},a^{1}_{i}]$. An action is chosen as $a={\rm arg\ max}_{a}Q(s,a)$.
III-B Simulation Platform
To bridge the gap between simulation and the real world, we propose a multi-agent training platform based on a physics-enhanced simulator such as Pybullet [29] and Gazebo. In most 2-D simulators, the actions of the agent are usually perfectly executed, and their positions are usually calculated by an equation like $x^{t+1}=x^{t}+v^{t}\Delta t+\frac{1}{2}a^{t}\Delta t^{2}$. In a physics-enhanced simulator, the agent state is computed by the physics engine, i.e., $x^{t+1}=\mathcal{T}(x^{t},v^{t},a^{t},f^{t},...)$, where $\mathcal{T}$ is the transfer function of the physics engine, $f^{t}$ is the force acting on the agent. The transfer function $\mathcal{T}$ makes the execution of the action randomness. A straightforward example is when the straight action executes in different types of simulators, the results differ. the 2-D simulator executes actions without deviation, whereas the agent in the physics-enhanced simulator diverges from the straight path, with even more pronounced deviations observed in real-world robots. This suggests that training in a physics-enhanced simulator enables the agent to learn more realistic policies, thereby improving its adaptability and facilitating seamless deployment on actual robots. To better train the policy, we add Gaussian noise to the agent’s perception to increase randomness further.
(a) 2-D simulator
(b) 3-D simulator
(c) Real robot
We propose a local replay mechanism for the trajectory collection of the agent’s collision, aiming to speed up the collection of multi-agent interaction trajectories. When the agent collides, the agent will be reset back to the state before the $N$ simulation steps so that the agent can quickly accumulate the collision data and speed up the training, where $N=300$ in this article. We propose this mechanism to address: (1) Spatiotemporal Sparsity: collisions in dense multi-agent scenarios are rare, making global reset methods inefficient; (2) Emergent Obstacle Prediction: sudden disturbances induce conservative behaviors (e.g., slowing/stopping), revealing implicit dynamic obstacle modeling; and (3) Asymmetric Adaptation: resetting single agents forces rapid adaptation to dynamic obstacles, simulating real-world pedestrian unpredictability and enhancing generalization.
Besides, to make the simulator generalize, we adopt the architecture of Gym [30], and we have diligently maintained the interfaces intending to provide users with a more convenient experience when utilizing the simulator. The simulator supports the random addition of random or queued agents or obstacles and controls the movement of each agent through an action set $\mathcal{A}$. At the end of a simulator step, the simulator returns an observed state set $\mathcal{O}$. The update frequency of the simulator can be modified to meet the policy training at different control frequencies. All of these functionalities are designed for easy reconfiguration. For example, to accommodate users seeking to evaluate Safe Reinforcement Learning (Safe-RL) [31], we have reconfigured our custom-designed simulator to include a constrained version by changing two functions of the origin version.
Figure 2: The overall framework diagram of our approaches is presented. The approaches can be divided into trajectory collection and policy update. During trajectory collection, each agent collects the trajectory independently and stores it in a replay buffer. After an action is executed, the simulation platform will return each agent’s reward and observation. After a specific number of iterations, the trajectories stored in the replay buffer will be used to calculate the loss, and the optimizer will then update the network based on the loss through gradient descent. The two stages alternate until a specific number of episodes is reached or the policy converges.
Algorithm 1 PPO For Multi-Agent Training
1:Initialize policy network $\pi_{\theta}$, value function network $V_{\phi}$, and set hyper-parameters
2:for $iteration=1,2,\ldots$ do
3:     for $t=1,2,\ldots,T_{\text{max}}$ do
4:         for each robot $i=1,2,\ldots,N$ do
5:              Run policy network $\pi_{\theta}$ to collect $\{o_{i}^{t},a_{i}^{t},r_{i}^{t}\}$
6:              Run value function network to collect $\{V_{\phi}(o_{i}^{t})\}$
7:         end for
8:     end for
9:     for each robot $i=1,2,\ldots,N$ do
10:         for $t=T_{\text{max}},T_{\text{max}}-1,\ldots,1$ do
11:              Calculate TD-error: $\delta_{i}^{t}=r_{i}^{t}+\gamma V_{\phi}(o_{i}^{t+1})-V_{\phi}(o_{i}^{t})$
12:              Estimate advantages using GAE: $A_{i}^{t}=\sum_{k=t}^{T_{\text{max}}}(\gamma\lambda)^{k-t}\delta_{i}^{k}$
13:              Calculate $V_{i}^{t}=A_{i}^{t}+V_{\phi}(o_{i}^{t})$
14:         end for
15:     end for
16:     Merge and shuffle the collected trajectories
17:     $\theta_{\text{old}}=\theta$, $\phi_{\text{old}}=\phi$
18:     for each epoch in Epoch do
19:         Calculate policy loss:
$L^{P}=-\mathbb{E}\left[\sum_{i=1}^{N}\sum_{t=1}^{T_{\text{max}}}\min\left(\frac{\pi_{\theta}(a_{i}^{t}|o_{i}^{t})}{\pi_{\theta_{\text{old}}}(a_{i}^{t}|o_{i}^{t})}A_{i}^{t},\text{clip}\left(\frac{\pi_{\theta}(a_{i}^{t}|o_{i}^{t})}{\pi_{\theta_{\text{old}}}(a_{i}^{t}|o_{i}^{t})},1-\epsilon,1+\epsilon\right)A_{i}^{t}\right)\right]$
20:         Calculate value loss:
$L^{V}=\max\left(\mathbb{E}\left[\sum_{i=1}^{N}\sum_{t=1}^{T_{\text{max}}}\left((V_{\phi}(o_{i}^{t})-V_{i}^{t})^{2},\left(\text{clip}(V_{\phi}(o_{i}^{t}),V_{\phi_{\text{old}}}(o_{i}^{t})-\epsilon,V_{\phi_{\text{old}}}(o_{i}^{t})+\epsilon)-V_{i}^{t}\right)^{2}\right)\right]\right)$
21:         Calculate entropy loss: $L^{\text{Entropy}}=\mathbb{E}[H(\pi_{\theta}(a_{i}^{t}|o_{i}^{t}))]$
22:         Calculate total loss: $L^{\text{Total}}=L^{P}+\alpha L^{V}+\beta L^{\text{Entropy}}$
23:         Update the parameters of $\pi_{\theta}$ and $V_{\phi}$ using Adam Optimizer with respect to $L^{\text{Total}}$
24:     end for
25:end for
III-C Training Process
We treat distributed multi-agent navigation as plenty of single agents navigating in an unknown environment, meaning each agent collects the trajectory respectively. As shown in Fig. 2, the training process can be divided into trajectory collection and policy update. During the trajectory collection, each agent first observes the LiDAR $o_{s}^{t}$, target point $o_{g}^{t}$, and velocity $o_{v}^{t}$. After that, it utilizes the policy and value network to calculate an action mean $[v^{t},\omega^{t}]$ and a value $v^{t}$. From these, an action $a^{t}$ is sampled based on the policy distribution $\pi_{\theta}(a^{t}|o^{t})$ and executed. The simulator then responds with a reward $r^{t}$, the next observable state $o^{t+1}$, and a termination signal $done^{t}$. The tuple of $<o^{t},a^{t},r^{t},o^{t+1},done^{t},V_{\phi}(o^{t})>$ is stored in replay buffer. This process is repeated until a specified number of iterations have been completed. After the trajectories are collected, the process proceeds to the policy update stage, where the trajectories stored in the replay buffer are used to calculate the total loss $L^{Total}$ for updating the policy and value networks. After updating the networks, the replay buffer will be cleared, and the agents collect trajectories again. The two stages alternate until convergence or episode limit is reached.
Algorithm 1 shows a more detailed training process. In lines 3-9, each agent collects the trajectories respectively and calculates the values using the value network under the same environment. After specific steps, in lines 10-15, according to the trajectories collected by each agent, the corresponding advantages $A$ of each trajectory are calculated through General Advantage Estimation (GAE) [32]. Then, we merge and shuffle the trajectories to get training data. In lines 17-24, we calculate the loss by:
$\begin{array}[]{l}L^{{P}}=-\text{min}(\frac{\pi_{\theta}(a|s)}{\pi_{\theta_{{old}}}(a|s)}A,\text{clip}(\frac{\pi_{\theta}(a|s)}{\pi_{\theta_{{old}}}(a|s)},1-\epsilon,1+\epsilon)A),\\ L^{V}=\max((V_{\theta}-V_{targ})^{2},(\operatorname{clip}(V_{\theta},V_{\theta_{old}}-\varepsilon,\\ \qquad\qquad\qquad\qquad\qquad\qquad\qquad V_{\theta_{old}}+\varepsilon)-V_{tart})^{2}),\\ L^{{Entropy}}=H(\pi_{\theta}(a^{t}|o^{t})),\end{array}$ (1)
where $V_{targ}=A+V_{\phi}(s)$. We get the total loss by weighted summing $L^{{Total}}=L^{{P}}+\alpha L^{V}+\beta L^{{Entopy}}$, where $\alpha$ and $\beta$ are hyperparameters. We use the Adam Optimizer [33] to update the policy and value network w.r.t $L^{Total}$. The algorithm proceeds to the next iteration until a specified iteration is reached. We also use some tricks to make the training more stable. The tricks include reward scaling [34], mini-batch updating, advantage normalizing, and learning rate decay.
We use two scenarios during training: single-agent and multi-agent. As we clip the distance input to the network, the size of the scenario does not affect the training. Thus, in the single-agent scenario, the initial point and the goal point of navigation are 10 m apart, and the multi-agent scenario’s initial and goal points are 16 m apart. The two scenarios are shown in Fig. 3.
(a) Single-agent scenario
(b) Multi-agent scenario
[…]
The measurement dimension $L_{n}$ of the LiDAR of the policy is different, and the max length of the LiDAR beam is 4 m. We first train the policy with $L_{n}=36$, which works successfully in the single-agent scenario. In the multi-agent scenario, the $L_{n}$ is updated to 72. Because if $L_{n}=36$, the agent may lose information when other agents stay in the internal of the LiDAR beam, with the biggest internal = 0.34 m, while $L_{n}=72$, the biggest internal = 0.17 m is enough for most situations. The field of view (FOV) is set to $\pi$ rad. Additionally, we stack the LiDAR observations of 5 consecutive frames to provide more comprehensive observation information. The agent’s velocity is set within the $[0,1]$ m/s, and the angular velocity is set within the $[-\pi,\pi]$ rad/s. The control frequency used in training is 60 Hz, which can make the reward variation of each action more apparent and maintain efficient control. Finally, the maximum number of simulation steps is set to 2500, sufficient for the agent to finish navigation through the optimal path.
III-D Reward Function
@@
$\begin{array}[]{l}R=r_{g}+r_{c}+r_{other},\\ r_{g}=\left\{\begin{array}[]{ll}r_{{arrival}}\qquad\qquad\quad\ \ \text{if }\left\|p_{i}^{t}-g_{i}\right\|<0.1,\\ w_{g}\left(\left\|p_{i}^{t-1}-g_{i}\right\|-\left\|p_{i}^{t}-g_{i}\right\|\right)\quad\text{otherwise,}\end{array}\right.\\ r_{c}=\left\{\begin{array}[]{ll}r_{{collision}}\qquad\qquad\text{if }d_{i}^{t}<0.01,\\ w_{c}(0.1-d_{i}^{t})\qquad\text{ if }d_{i}^{t}<0.1,\\ 0\qquad\qquad\qquad\quad\text{ otherwise,}\end{array}\right.\\ \end{array}$ (2)
where $r_{other}$ is the reward that varies in different approaches, $r_{g}$ is the distance reward, which is computed as the difference in the distance between the robot and the goal point at the current time step and the previous time step, $r_{c}$ is the obstacle penalty, $w_{g}$ and $w_{c}$ are the weight of $r_{g}$ and $r_{c}$ respectively, $d_{i}^{t}$ means the smallest distance between agent $i$ and obstacles at time step $t$. However, this reward function setup may make agents learn conservative policy. Additionally, many works impose control constraints, such as velocity and angular velocity, on $r_{other}$ to maintain action coherence. Too many constraints may lead to instability in the training process and difficulty adjusting the weights between various terms. Moreover, in some scenarios, the reward function does not adequately indicate optimal actions.
@@
Figure 4: Scenarios where the classical reward function demonstrates significant limitations.
As shown in Fig. 4, when an agent needs to navigate through the gap between two obstacles, it may incur penalties due to $r_{c}$ when approaching them, even though traversing the gap is the optimal action. As a result, the agent may circumnavigate obstacles. Furthermore, agents focus exclusively on proximal threats in multi-agent scenarios while ignoring potentially hazardous agents at greater distances, resulting in unsafe navigation decisions. To address the issues above, we have introduced modifications to the $r_{c}$ component, specifically by:
$\begin{array}[]{l}r^{\prime}_{c}=\left\{\begin{array}[]{ll}r_{collision}&\text{if }d_{i}^{t}<0.01,\\ w^{\prime}_{c}\left(l-sl_{i}^{t}\right)+r_{c}&\text{ otherwise,}\end{array}\right.\end{array}$ (3)
where $r_{arrival}=5$, $r_{collision}=-5$, $w_{g}=5$, $w^{\prime}_{c}=-0.25$, $w_{c}=-0.03$, $l$ represents the maximum length of the LiDAR rays, $sl_{i}^{t}$ represents the LiDAR observations of agent $i$ at time step $t$, $w^{\prime}_{c}$ is a weight vector that varies with angular velocity, with the weight decreasing in $\frac{1}{\sqrt{2\pi}\sigma}e^{-\frac{(x-\omega\Delta t)^{2}}{2\sigma^{2}}}$ fashion from $\omega\Delta t$ towards both sides, where $\sigma=0.2$, $\Delta t$ is the time interval of the control. The schematic diagram of $w^{\prime}_{c}$ is shown in Fig. 5.
@@
Using $r^{\prime}_{c}$, the policy pays more attention to the direction of the current angular velocity while maintaining the closest distance to the surrounding obstacles. The reward based on angular velocity of the agent allows the policy to better cope with the mentioned situations and enable the agent to pay more attention to the choice of angular velocity.
III-E Network Structure
The agent only requires more precise, goal-oriented actions when nearing the target location, whereas general directional movement suffices otherwise. We clip the agent-goal distance rather than feeding raw distance values into the network to enable the policy to differentiate between these two regimes explicitly. This design decouples the policy’s decision-making from absolute distance metrics, ensuring scalability to larger environments while maintaining consistent behavioral logic across spatial scales.
To address the loss of information in POMDP, we first employ LSTM Layers to encode LiDAR observations. By stacking sequential LiDAR observations, the LSTM module captures historical observation memory and integrates consecutive observation patterns. While the policy performs well in static environments and 6-agent test scenarios, its performance degrades significantly when the number of agents exceeds 6. This limitation arises because the LSTM-based policy overly focuses on LiDAR observations while neglecting the agent’s velocity, making it difficult to navigate effectively in denser multi-agent environments.
To address this issue, we introduce an attention mechanism that dynamically balances the importance of LiDAR observations and velocity information. First, the velocity and goal information are encoded through fully connected (FC) layers to produce the query vector $Q$, while the LiDAR observations are processed by the LSTM module with FC layers to generate the value $V$ and key $K$ vectors, respectively. The attention mechanism then integrates all relevant information into a comprehensive representation, as described in (4). Integrating the proposed reward function enhances policy and achieves robust navigation in complex multi-agent scenarios. The architecture, which combines LSTM with an attention mechanism, is designated the LSTM-Attention policy.
TABLE I: Experiment with 1 robot and 15 obstacles.
success rate(%) collision rate(%) trap rate (%) average step
IJRR [12] 91.3 8.6 0.1 1383.23
  Linear policy 67.2 2.4 30.4 550.40
LSTM policy 93.1 6.9 0.0 1187.73
LSTM-Attention 95.9 3.9 0.2 1300.78
  (Safe) LSTM-Attention 97.6 2 0.04 1353.63
TABLE II: Experiment with 6 robots and 5 obstacles.
success rate(%) collision rate(%) trap rate (%) average step
NH-ORCA [6] 98.4 0.33 1.27 1198.82
IJRR [12] 85.03 14.63 0.33 1617.84
LSTM policy 91.89 8.15 0.0 981.01
LSTM-Attention 94.78 4.68 0.53 1805.94